# An oscillating Rastall universe crossing the phantom divide line Nasr Ahmed 1,2, Anirudh Pradhan3, Archana Dixit 4

### v1

@@ 1 Introduction and Motivation
Bouncing scenarios (also named as cyclic cosmology) have been firstly suggested as an alternative way to solve some of the problems of the standard model of cosmology such as flatness and initial singularity. The universe in the singularity-free Big Bounce emerges from a previous contracting stage [16, 17, 18, 19] (see [20] for a review) where the contraction-expansion cycle is supposed to continue forever. There have been an extensive interest in cyclic cosmological models in the literature. they have been studied in many modified gravity theories [18, 21, 22, 23, 24, 25, 19, 26]. A general cyclic model in $f(R)$ gravity has been constructed in [27]. A unitary version of Conformal Cyclic Cosmology ( CCC ) has been suggested in [28]. An M-theory inspired cyclic Model with two branes has been presented in [29] where the EoS parameter satisfies $\omega\gg 1$ during the contraction phase. A new novel cyclic theory of the universe has been presented in [30] where the scale factor grows exponentially from one cycle to another. This recent cyclic model resolves many problems and produces a nearly scale invariant spectrum of density perturbations. An interesting singularity-free oscillatory model has been discussed in [31] where the late-time evolution was probed in the context of Quasi Steady State cosmology QSSC [32, 33, 34] to alleviate the Hubble tension. The cosmic scale factor parametrization has been chosen according to the description in [32, 33, 34, 35, 36] as
@@
Rastall gravity is one such modification that was proposed by Peter Rastall in 1972 [37]. Due to its presentation as a modified theory of gravity with a non-conserved stress energy tensor and a peculiar non-minimal coupling between geometry and matter, Rastall gravity (RG) has emerged as a leading theory of gravity in modern times. The conservation law that demonstrated a divergence free energy momentum tensor in general relativity (GR) has been accepted; that is, $\nabla_{\nu}T_{\mu\nu}=0$, where $\nabla_{\nu}$ is the covariant derivative. Rastall, on the other hand, adopted a new and distinct conservation law that conjectured the GR concepts. In this theory, the energy momentum tensor does not a conserved quantity. Rastall theorem is defined as $\nabla_{\mu}=\left(\frac{\kappa}{16\pi}\right)\nabla_{\mu}R$, where $\kappa$ is the coupling constant and $R$ be the Ricci scalar. The second Bianchi identity in this theory remains the same i.e., $\nabla^{\mu}G_{\nu\mu}=0$, where $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}$ be the Einstein tensor. One significant aspect of RG is that, while the mathematical portion of the theory remains invariant, modifications have only been made to the matter portion [38, 39, 40]. Das et al. [41] demonstrated that RG is equal to Einstein gravity and offered some cosmological implications in the structure of modified RG. Holographic dark energy in RG was introduced by Ghaffari et al. [42], taking into account vacuum energy, which serves as DE. They accepted the idea that the current accelerating cosmos is caused by the sum of this energy and the Rastall term. With the assumption that the universe is made up of interacting/non-interacting dark energy (DE) and dark matter (DM), Saleem and Shahnila [43] investigated the phenomenon of cosmic evolution using curved FLRW space-time bounded by an apparent horizon with a particular holographic cut-off in the framework of Rastall gravity. Saleem et al. [44] recently introduced constant-roll warm inflation, a novel method for determining the precise inflationary solutions to the Friedman equations within the framework of Rastall theory of gravity (RTG). Using a linear parametrization of the Equation of State (EoS) in FLRW background, Singh et al. [45] examined the evolution of the universe within the framework of RTG. In the framework of RTG, Saleem and Hassan [46] studied the dynamics of warm inflation induced by vector fields and concluded that the modified theory was compatible with the 2018 Planck observational data. In various contexts, several researchers have examined the cosmological scenario within the framework of Rastall gravity [47, 48, 49, 50, 51, 52, 53, 54, 55].

### v2

@@ 1 Introduction and Motivation
Bouncing scenarios (also named as cyclic cosmology) have been firstly suggested as an alternative way to solve some of the problems of the standard model of cosmology such as flatness and initial singularity. The universe in the singularity-free Big Bounce emerges from a previous contracting stage [16, 17, 18, 19] (see [20] for a review) where the contraction-expansion cycle is supposed to continue forever. There have been an extensive interest in cyclic cosmological models in the literature. they have been studied in many modified gravity theories [18, 19, 21, 22, 23, 24, 25, 26, 27]. A general cyclic model in $f(R)$ gravity has been constructed in [28]. A unitary version of Conformal Cyclic Cosmology ( CCC ) has been suggested in [29]. An M-theory inspired cyclic Model with two branes has been presented in [30] where the EoS parameter satisfies $\omega\gg 1$ during the contraction phase. A new novel cyclic theory of the universe has been presented in [31] where the scale factor grows exponentially from one cycle to another. This recent cyclic model resolves many problems and produces a nearly scale invariant spectrum of density perturbations. An interesting singularity-free oscillatory model has been discussed in [32] where the late-time evolution was probed in the context of Quasi Steady State cosmology QSSC [33, 34, 35] to alleviate the Hubble tension. The cosmic scale factor parametrization has been chosen according to the description in [33, 34, 35, 36, 37] as
@@
Rastall gravity is one such modification that was proposed by Peter Rastall in 1972 [38]. Due to its presentation as a modified theory of gravity with a non-conserved stress energy tensor and a peculiar non-minimal coupling between geometry and matter, Rastall gravity (RG) has emerged as an interesting and important theory of gravity in modern times. The conservation law that demonstrated a divergence free energy momentum tensor in general relativity (GR) has been accepted; that is, $\nabla_{\nu}T^{\mu\nu}=0$, where $\nabla_{\nu}$ is the covariant derivative. Rastall, on the other hand, adopted a new and distinct conservation law that conjectured the GR concepts. In this theory, the energy momentum tensor does not a conserved quantity. Rastall formulation is defined as $\nabla_{\mu}T^{\mu}_{\nu}=\left(\frac{\kappa}{16\pi}\right)\nabla_{\nu}R$, where $\kappa$ is the coupling constant and $R$ be the Ricci scalar. The second Bianchi identity in this theory remains the same i.e., $\nabla^{\mu}G_{\nu\mu}=0$, where $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}$ be the Einstein tensor. One significant aspect of RG is that, while the mathematical portion of the theory remains invariant, modifications have only been made to the matter portion [39, 40, 41]. For the Rastall gravity framework, Visser [42] have claimed that the Rastall gravity may have the complete equivalence with the General relativity. However, Darabi et al [43] have discussed that the Rastall gravity may be visualized as a modified gravity theory and it is different from the General relativity. Das [44] demonstrated that RG is equal to Einstein gravity and offered some cosmological implications in the structure of modified RG. Recently, Golovnev [45] have claimed that Rastall gravity may be equivalent to the General relativity by analyzing the claims of Visser [42] as well as Darabi et al [43]. The derivation of action for Rastall gravity have been given subjected to the condition $\sqrt{-g}=1$ [45]. Holographic dark energy in RG was introduced by Ghaffari et al. [46], taking into account vacuum energy, which serves as DE. They accepted the idea that the current accelerating cosmos is caused by the sum of this energy and the Rastall term. With the assumption that the universe is made up of interacting/non-interacting dark energy (DE) and dark matter (DM), Saleem and Shahnila [47] investigated the phenomenon of cosmic evolution using curved FLRW space-time bounded by an apparent horizon with a particular holographic cut-off in the framework of Rastall gravity. Saleem et al. [48] recently introduced constant-roll warm inflation, a novel method for determining the precise inflationary solutions to the Friedman equations within the framework of Rastall theory of gravity (RTG). Using a linear parametrization of the Equation of State (EoS) in FLRW background, Singh et al. [49] examined the evolution of the universe within the framework of RTG. In the framework of RTG, Saleem and Hassan [50] studied the dynamics of warm inflation induced by vector fields and concluded that the modified theory was compatible with the 2018 Planck observational data. In various contexts, several researchers have examined the cosmological scenario within the framework of Rastall gravity [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. This theory have received a widespread attention to discuss and explain different phenomenon related to the universe evolution. As a fact, we aim to study the universe evolution exhibiting oscillatory expansion with analysis of thermodynamic behavior. It is worthwhile to mention that Rastall gravity possess oscillatory solutions subjected to the fixed point which are marginally stable [57, 58].